# Non-Convex Stochastic Composite Optimization with Polyak Momentum

### v1

@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is smooth and $\psi\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is not necessarily differentiable, but a simple function. Such paradigm is ubiquitous in machine learning and beyond, and it covers a wide range of generalizations of the vanilla optimization problem, including regularized machine learning problems [26], signal processing [6], image processing [28] and many more. It also naturally covers constrained optimization problems by considering the indicator function of the constraint set, and some recent variants of such ideas have been applied to distributed and federated machine learning problems [29].
The problem formulation (1) is often used in Machine Learning as the proximal term $\phi$ allows encoding prior domain knowledge. For instance, it can impose constraints on the variables, handle non-differentiability, induce sparsity, or it can take the form of a regularizer. Recently, there has been a surge of interest in these preconditioning techniques in the Deep Learning context [17, 41]. Given a loss objective function $\ell\colon\mathbb{R}^{d}\to\mathbb{R}$, an auxiliary function (or preconditioner) $\hat{\ell}$ is constructed—for instance, by approximating $\ell$ with a subset of the data samples or synthetic samples. This results in the formulation:
@@
This is a special case of Problem (1), with $f\coloneqq\ell-\hat{\ell}$ and $\psi\coloneqq\hat{\ell}$. This formulation has the advantage that the training can be accelerated if the proxy function $\hat{\ell}$ remains simple to optimize. These applications motivate the need to further our understanding of the stochastic composite optimization problem in the non-convex regime, especially where large batches are less preferred or even unavailable [24, 34].
In the convex and strongly convex case, the complexity of solving Problem (1) is well understood [16, 14]. On the other hand, the existing theory in the non-convex regime is unsatisfactory. In particular, when the gradient noise has variance $\sigma^{2}$, with the vanilla algorithm, the squared norm of the gradient of $F$ is only shown to converge to $\mathcal{O}(\nicefrac{{1}}{{K}})+\Omega(\sigma^{2})$ after $K$ iterations. This implies that to converge to an arbitrary $\varepsilon$ error, one needs to take mega batches of size $\Omega(\varepsilon^{-1})$ at each iteration of the algorithm, and the total number of stochastic gradient oracle calls is $\Omega(\varepsilon^{-2})$. In practice and theory, smaller-batch methods are often preferred over mega-batch methods [^1] By a mega-batch we mean a batch of the size of the order $\Omega(\varepsilon^{-1})$.. For example, mega-batch methods follow the full gradient methods much more closely at each step than smaller batch methods, and it is observed empirically to be adversarial to the generalization performance [40, 24, 20], and theoretically [36] for certain Machine Learning tasks. Furthermore, in many practical settings, mega-batches are unavailable or intractable to sample, e.g., in medical tasks [34]; federated Reinforcement Learning [21, 19]; and multi-agent Reinforcement Learning [10].
In this work, we revisit the stochastic composite optimization problem in the non-convex regime and show that the Polyak momentum technique addresses the aforementioned mega-batch issue while retaining the optimal convergence rate $\mathcal{O}(\varepsilon^{-2})$. Polyak momentum is the de-facto standard for training Deep Learning models in practice [22], and understanding its theoretical efficacy in various settings is an active research direction [7, 12].
@@ 1.1 Stochastic proximal gradient method
To solve the stochastic composite optimization problem (1), we consider the stochastic proximal gradient method, which solves the following subproblem in each iteration (commonly known as the proximal step):
$\mathbf{x}_{k+1}$ $\coloneqq\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{d}}\{\Omega_{k}(\mathbf{x})\}\,,\text{with}$ (2)
$\Omega_{k}(\mathbf{x})$ $\coloneqq\langle\mathbf{g}_{k},\mathbf{x}\rangle+\psi(\mathbf{x})+\frac{M}{2}\left\lVert\mathbf{x}-\mathbf{x}_{k}\right\rVert^{2}\,.$
Here, $\mathbf{g}_{k}$ denotes a random vector computed at each iteration $k$ using the stochastic first-order oracle of $f$, and $M>0$ is a regularization (or stepsize) parameter. For the vanilla stochastic proximal gradient method, $\mathbf{g}_{k}$ is a stochastic gradient of $f$ at $\mathbf{x}_{k}$. We note that when $\psi\equiv 0$, this reduces to the vanilla SGD with step size $\frac{1}{M}$. When $\psi\not\equiv 0$, it has long been known that if error-dependent batches are not allowed, then such a method is only proved to converge to a neighborhood of the stationary point of $F$ up to the variance of the gradient noise [16]. In Section 4 we give a simple example to illustrate this phenomenon. Therefore, one needs $\varepsilon$-dependent mini-batches (or mega-batches) to converge to an $\varepsilon$ neighborhood of the stationary point of $F$.
@@ 1.2 Our contributions
Stochastic Polyak momentum has seen a huge success in practice [22]. In this work, we study the effect of Polyak momentum in the non-convex regime and provide a theoretical analysis of momentum’s stabilizing effect in the stochastic composite optimization setting, particularly in the small batch regime.
@@ 2 Related works
There is a huge body of work on stochastic and composite optimization. Here, we focus on the stochastic and non-convex regime and do not get into the details of the works in convex (e.g. [13, 15]) or deterministic regime (see [31]). The first work that considers the stochastic composite optimization problem in the non-convex case appears to be [16], in which the convergence of the stochastic proximal gradient method to a neighborhood of the stationary point of $F$ up to the variance of the gradient noise is established. If mega-batches are allowed, the paper shows that the algorithm takes asymptotically $\mathcal{O}(\varepsilon^{-2})$ stochastic gradient oracle calls to converge to an $\varepsilon$ neighborhood of the stationary point of $F$. These results were later extended to incorporate the acceleration technique, but the convergence still requires mega-batches. In particular, the method requires $k$ samples at the $k$th iteration while the total number of oracle calls is not improved [14]. Even though these methods require mega-batches, their upper bounds on the total number of stochastic gradient oracle calls match the lower bound asympotically [2].
Average smoothness and variance reduction
In another research direction, to break the $\mathcal{O}(\varepsilon^{-2})$ lower bound for stochastic non-convex optimization, many variance reduction techniques have been proposed, at the cost of two additional assumptions on the objective and problem structure:
  • $f$ is $\widetilde{L}$-average smooth: ${\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x},\xi)-\nabla f(\mathbf{y},\xi)\right\rVert^{2}\right]\leq\widetilde{L}^{2}\left\lVert\mathbf{x}-\mathbf{y}\right\rVert^{2}$. Note that this assumption might not be satisfied for some simple smooth functions, and it is strictly stronger than the standard smoothness assumption. [^2] By Jensen’s inequality, we have that any $\widetilde{L}$-average smooth function is also $\widetilde{L}$-smooth, see also 3.3. There exists $L$-smooth functions that are not average smooth.
  • These variance reduction techniques require the access to two stochastic gradients $\nabla f(\mathbf{x}_{k},\xi_{k})$ and $\nabla F(\mathbf{x}_{k-1},\xi_{k})$ at iteration $k$, which might not always be available in practice (e.g. [10, 5]).
There have been a long line of works on variance reduction techniques, including Prox-Spiderboost [39] (composite variant of Spider [11]), Hybrid-SGD [37], and PStorm [42] (composite variant of Storm [8]). These methods achieve a $\mathcal{O}(\varepsilon^{-\nicefrac{{3}}{{2}}})$ asymptotic complexity.
@@
The first-order convergence criteria for non-convex optimization problems is an ongoing research topic. In most of the works studying the non-convex composite optimization problem (1), the convergence is studied in terms of the proximal gradient mapping [16, 14, 39, 37, 42]. There have been several works proposing different convergence criteria for non-convex optimization problems and discussing how these definitions affect the corresponding convergence complexities, e.g. [9, 44, 45], but these are parallel to our work, and we do not discuss them in detail here. Our work mostly focuses on the convergence in terms of $\left\lVert\nabla F(\mathbf{x}_{k})\right\rVert^{2}$. We note that all out theory works with proximal gradient mapping criteria, see Appendix D.
@@
The idea of using Polyak momentum (a.k.a. heavy-ball momentum) was first proposed in [32] for strongly convex quadratic objective in the deterministic case. The first non-asymptotic analysis of SGD with Polyak momentum in the smooth non-convex regime is given in [43] and was later refined in [27]. Some recent works also studied how Polyak momentum can be used to remove the dependence on large batches in various settings, e.g. for normalized SGD [7] and for communication compressed optimization [12]. The analysis of stochastic gradient methods with Polyak momentum remains an active research topic [38, 35, 25, 18].
@@ 3 Problem formulation and assumptions
We study the convergence of the algorithms in terms of $\left\lVert\nabla F(\mathbf{x})\right\rVert^{2}$. When $\psi$ is convex and non-differentiable, $\nabla F$ is understood as an appropriate choice of the subgradient. We do not consider the non-convex and non-differentiable case for $\psi$, as this is an exotic scenario in the literature. For simplicity, we assume differentiability when presenting the proofs in the main text, but all our results hold for the convex and non-differentiable case for $\psi$.
Note that most of the existing works on stochastic composite optimization assume that $\psi$ is convex and analyze the convergence of their algorithms in terms of the proximal gradient mapping [16, 14, 39, 37, 42], which is closely related to what we consider in this work. In Appendix D, we discuss the proximal gradient mapping and argue that our definition is more natural in the non-convex regime.
@@
Assumption 3.1.
We assume that, at each iteration $k$ of the algorithms, the output $\mathbf{x}_{k+1}$ of the proximal step (2) satisfies the following:
$\Omega_{k}(\mathbf{x}_{k+1})\leq\Omega_{k}(\mathbf{x}_{k})\,.$
@@
Now, we introduce the smoothness assumption on $f$. Note that this is a weaker assumption than the average smoothness assumption used in the variance reduction type method [39, 37, 42].
Assumption 3.3.
@@
This is a standard assumption in the analysis of stochastic gradient methods [30, 4]. Taking mini-batches is equivalent to dividing the variance by the batch size.
@@ 4 Lower bound for the vanilla stochastic proximal gradient method
Proposition 4.1.
For any step-size coefficient $M>0$, there exists $F=f+\psi:\mathbb{R}^{d}\to\mathbb{R}$ with $L$-smooth $f$ and stochastic gradient oracle of $f$ with variance $\sigma^{2}$, such that the method (2) with $\mathbf{g}_{k}$ being the stochastic gradient of $f$ at $\mathbf{x}_{k}$ cannot converge to the stationary point of $F$ beyond a $\frac{\sigma^{2}}{4}$ neighborhood. In particular, given any $\mathbf{x}_{k}$, ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\geq\frac{\sigma^{2}}{4}$.
@@
Consider $f(\mathbf{x})=\frac{L\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ and $\psi(\mathbf{x})=\frac{a\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ for some $a>0$. The stochastic gradient oracle of $f$ is $\mathbf{g}_{k}=\nabla f(\mathbf{x}_{k})+\xi_{k}$ where $\xi_{k}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. Clearly, the construction satisfies 3.3 and 3.4. It is easy to check that the proximal step is given by:
$\mathbf{x}_{k+1}=\frac{M-L}{M+a}\mathbf{x}_{k}-\frac{\xi_{k}}{M+a}\,.$
Therefore,
${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $=\frac{(M+L)^{2}}{(M+a)^{2}}{\mathbb{E}}\left[\left\lVert(M-L)\mathbf{x}_{k}-\xi_{k}\right\rVert^{2}\right]$
$=\frac{(M+L)^{2}}{(M+a)^{2}}\left((M-L)^{2}\left\lVert\mathbf{x}_{k}\right\rVert^{2}+\sigma^{2}\right)$
$\geq\frac{(M+L)^{2}\sigma^{2}}{(M+a)^{2}}\,,$
where the first equality follows from the independence of $\xi_{i}$ and $\mathbf{x}_{k}$. Hence, for any $a>M$, we must have that method (2) with the stochastic gradient $\mathbf{g}_{k}$ cannot converge to the stationary point of $F$ beyond $\frac{\sigma^{2}}{4}$ neighborhood. ∎
Remark 4.2.
We point out that one of the characteristics of the proximal gradient method is that it is oblivious to the choice of $\psi$. In particular, from a theoretical point of view, the algorithm is only aware of the parameters of $f$, and the convergence guarantees should always hold despite of the choice of $\psi$.
The problem is that the variance of the gradient noise keeps the iterates away from the stationary point of $F$, and we need some mechanism to reduce the variance of the gradient noise towards the end of the optimization process. This is when the Polyak momentum comes into play.

### v4

@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is smooth and $\psi\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is not necessarily differentiable, but a simple function. Such paradigm is ubiquitous in machine learning and beyond, and it covers a wide range of generalizations of the vanilla optimization problem, including regularized machine learning problems Liu et al. (2015), signal processing Combettes and Pesquet (2010), image processing Luke (2020) and many more. It also naturally covers constrained optimization problems by considering the indicator function of the constraint set, and some recent variants of such ideas have been applied to distributed and federated machine learning problems Mishchenko et al. (2022).
The problem formulation (1) is often used in Machine Learning as the proximal term $\psi$ allows encoding prior domain knowledge. For instance, it can impose constraints on the variables, handle non-differentiability, induce sparsity, or it can take the form of a regularizer. Recently, there has been a surge of interest in these preconditioning techniques in the Deep Learning context Hendrikx et al. (2020); Woodworth et al. (2023). Given a loss objective function $\ell\colon\mathbb{R}^{d}\to\mathbb{R}$, an auxiliary function (or preconditioner) $\hat{\ell}$ is constructed—for instance, by approximating $\ell$ with a subset of the data samples or synthetic samples. This results in the formulation:
@@
This is a special case of Problem (1), with $f\coloneqq\ell-\hat{\ell}$ and $\psi\coloneqq\hat{\ell}$. This formulation has the advantage that the training can be accelerated if the proxy function $\hat{\ell}$ remains simple to optimize. These applications motivate the need to further our understanding of the stochastic composite optimization problem in the non-convex regime, especially where large batches are less preferred or even unavailable LeCun et al. (2012); Rieke et al. (2020).
In the convex and strongly convex case, the complexity of solving Problem (1) is well understood Ghadimi et al. (2016); Ghadimi and Lan (2013a). On the other hand, the existing theory in the non-convex regime is unsatisfactory. In particular, when the gradient noise has variance $\sigma^{2}$, with the vanilla algorithm, the squared norm of the gradient of $F$ is only shown to converge to $\mathcal{O}(\nicefrac{{1}}{{K}})+\Omega(\sigma^{2})$ after $K$ iterations. This implies that, to converge to an arbitrary $\varepsilon$ error, one needs to take mega batches of size $\Omega(\varepsilon^{-1})$ at each iteration of the algorithm to reduce the stochastic noise term; and the total number of stochastic gradient oracle calls is $\Omega(\varepsilon^{-2})$. In practice and theory, smaller-batch methods are often preferred over mega-batch methods. For example, mega-batch methods follow the full gradient methods much more closely at each step than smaller batch methods, and it is observed empirically to be adversarial to the generalization performance Wilson and Martinez (2003); LeCun et al. (2012); Keskar et al. (2017), and theoretically Sekhari et al. (2021) for certain Machine Learning tasks. Furthermore, in many practical settings, mega-batches are unavailable or intractable to sample, e.g., in medical tasks Rieke et al. (2020); federated Reinforcement Learning Khodadadian et al. (2022); Jin et al. (2022); and multi-agent Reinforcement Learning Doan et al. (2019).
In this work, we revisit the stochastic composite optimization problem in the non-convex regime and show that the Polyak momentum technique addresses the aforementioned mega-batch issue while retaining the optimal convergence rate $\mathcal{O}(\varepsilon^{-2})$. Polyak momentum is the de-facto standard for training Deep Learning models in practice Kingma and Ba (2014), and understanding its theoretical efficacy in various settings is an active research direction Cutkosky and Mehta (2020); Fatkhullin et al. (2023).
@@ 1.1 Stochastic Proximal Gradient Method
To solve the stochastic composite optimization problem (1), we consider the stochastic proximal gradient method, which solves the following subproblem in each iteration (commonly known as the proximal step): find $\mathbf{x}_{k+1}$ such that
$\Omega_{k}(\mathbf{x}_{k+1})\leq\Omega_{k}(\mathbf{x}_{k})\quad\text{and}\quad\nabla\Omega_{k}(\mathbf{x}_{k+1})=0,$
where $\Omega_{k}(\mathbf{x})\coloneqq\langle\mathbf{g}_{k},\mathbf{x}\rangle+\psi(\mathbf{x})+\frac{M_{k}}{2}\left\lVert\mathbf{x}-\mathbf{x}_{k}\right\rVert^{2}$. Here, $\mathbf{g}_{k}$ denotes a random vector computed at each iteration $k$ using the stochastic first-order oracle of $f$, and $M_{k}>0$ is a regularization (or stepsize) parameter. $M_{k}$ can be a user defined constant, or it can be a non-decreasing sequence. For the vanilla stochastic proximal gradient method, $\mathbf{g}_{k}$ is a stochastic gradient of $f$ at $\mathbf{x}_{k}$.
When $\psi$ is convex, the proximal step is equivalent to minimizing $\Omega_{k}$:
$\mathbf{x}_{k+1}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{d}}\Bigl{\{}\langle\mathbf{g}_{k},\mathbf{x}\rangle+\psi(\mathbf{x})+\frac{M_{k}}{2}\left\lVert\mathbf{x}-\mathbf{x}_{k}\right\rVert^{2}\Bigr{\}}.$ (2)
We note that when $\psi\equiv 0$, this reduces to the vanilla SGD with step size $\frac{1}{M_{k}}$. When $\psi\not\equiv 0$, it has long been known that if error-dependent batch sizes are not allowed, then such a method is only proved to converge to a neighborhood of the stationary point of $F$ up to the variance of the gradient noise Ghadimi et al. (2016). In Section 4 we give a simple example to illustrate this phenomenon. Therefore, one needs $\varepsilon$-dependent mini-batches (or mega-batches) to converge to an $\varepsilon$ neighborhood of the stationary point of $F$.
@@ 1.2 Our Contributions
Stochastic Polyak momentum has seen a huge success in practice Kingma and Ba (2014). In this work, we study the effect of Polyak momentum in the non-convex regime and provide a theoretical analysis of momentum’s stabilizing effect in the stochastic composite optimization setting, particularly in the small batch regime.
@@ 2 Related Works
There is a huge body of work on stochastic and composite optimization. Here, we focus on the stochastic and non-convex regime and do not get into the details of the works in the convex (e.g. Ghadimi and Lan, 2012, 2013b) or the deterministic regime (see Nesterov, 2013). The first work that considers the stochastic composite optimization problem in the non-convex case appears to be Ghadimi et al. (2016), in which they established the convergence of the stochastic proximal gradient method to a neighborhood of the stationary point of $F$ up to the variance of the gradient noise. If mega-batches are allowed, they showed that the algorithm takes asymptotically $\mathcal{O}(\varepsilon^{-2})$ stochastic gradient oracle calls to converge to an $\varepsilon$ neighborhood of the stationary point of $F$. It was later extended to incorporate the acceleration technique, but the convergence still requires mega-batches. In particular, it requires $k$ samples at the $k$th iteration while the total number of oracle calls is not improved Ghadimi and Lan (2013a). Even though these methods require mega-batches, their upper bounds on the total number of stochastic gradient oracle calls match the lower bound asympotically Arjevani et al. (2023).
In another research direction, to break the $\mathcal{O}(\varepsilon^{-2})$ lower bound for stochastic non-convex optimization, there have been a long line of works on variance reduction techniques, including Prox-Spiderboost Wang et al. (2019) (composite variant of Spider Fang et al. (2018)), Hybrid-SGD Tran-Dinh et al. (2022), and PStorm Xu and Xu (2022) (composite variant of Storm Cutkosky and Orabona (2019)). These methods achieve a $\mathcal{O}(\varepsilon^{-\nicefrac{{3}}{{2}}})$ asymptotic complexity at the cost of two additional assumptions on the objective and problem structure:
  • $f$ is $\widetilde{L}$-average smooth: ${\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x},\xi)-\nabla f(\mathbf{y},\xi)\right\rVert^{2}\right]\leq\widetilde{L}^{2}\left\lVert\mathbf{x}-\mathbf{y}\right\rVert^{2}$. Note that this assumption might not be satisfied for some simple smooth functions, and it is strictly stronger than the standard smoothness assumption. [^1] By Jensen’s inequality, we have that any $\widetilde{L}$-average smooth function is also $\widetilde{L}$-smooth, see also Assumption 3. There exists $L$-smooth functions that are not average smooth.
  • These variance reduction techniques require the access to two stochastic gradients $\nabla f(\mathbf{x}_{k},\xi_{k})$ and $\nabla F(\mathbf{x}_{k-1},\xi_{k})$ at iteration $k$, which might not always be available in practice (e.g. Doan et al., 2019; Chen et al., 2022).
@@
The first-order convergence criteria for non-convex optimization problems is an ongoing research topic. Our work mostly focuses on the convergence in terms of $\left\lVert\nabla F(\mathbf{x}_{k})\right\rVert^{2}$, which we think is the most natural criteria. Many early works study the convergence of problem (1) in terms of the proximal gradient mapping Ghadimi et al. (2016); Ghadimi and Lan (2013a); Wang et al. (2019); Tran-Dinh et al. (2022); Xu and Xu (2022). More recently, there have been several works proposing to study the convergence in terms of Moreau envelope and discussing how the different criterias affect the corresponding convergence complexities Davis and Drusvyatskiy (2019); Zhang et al. (2020). We discuss the differences and connections between these definitions in Appendix G.
@@
The idea of using Polyak momentum (a.k.a. heavy-ball momentum) was first proposed in Polyak (1964) for strongly convex quadratic objective in the deterministic case. The first non-asymptotic analysis of SGD with Polyak momentum in the smooth non-convex regime is given in Yu et al. (2019) and was later refined in Liu et al. (2020). Some recent works also studied how Polyak momentum can be used to remove the dependence on large batches in various settings, e.g. for normalized SGD Cutkosky and Mehta (2020) and for communication compressed optimization Fatkhullin et al. (2023). The analysis of stochastic gradient methods with Polyak momentum remains an active research topic Wang and Abernethy (2020); Sebbouh et al. (2021); Li et al. (2022); Jelassi and Li (2022).
@@ 3 Problem Formulation and Assumptions
In the main body of the paper we study the convergence of the algorithms in terms of $\left\lVert\nabla F(\mathbf{x})\right\rVert^{2}$ where we assume that $\psi$ is differentiable. We discuss the case where $\psi$ is convex and non-differentiable in Appendix F. We do not consider the non-convex and non-differentiable case for $\psi$, as this is an exotic scenario in the literature.
Note that most of the existing works on stochastic composite optimization assume that $\psi$ is convex and analyze the convergence of their algorithms in terms of the proximal gradient mapping Ghadimi et al. (2016); Ghadimi and Lan (2013a); Wang et al. (2019); Tran-Dinh et al. (2022); Xu and Xu (2022), which is closely related to what we consider in this work. In Appendix G, we discuss the proximal gradient mapping and argue that our definition is more natural in the non-convex regime.
@@
Assumption 1.
We assume that, at each iteration $k$ of the algorithms, the output $\mathbf{x}_{k+1}$ of the proximal step satisfies $\Omega_{k}(\mathbf{x}_{k+1})\leq\Omega_{k}(\mathbf{x}_{k})$.
@@
Now, we introduce the smoothness assumption on $f$. Note that this is a weaker assumption than the average smoothness assumption used in the variance reduction type methods Wang et al. (2019); Tran-Dinh et al. (2022); Xu and Xu (2022).
Assumption 3.
@@
This is a standard assumption in the analysis of stochastic gradient methods Nemirovskij and Yudin (1983); Bubeck et al. (2015). Taking mini-batches is equivalent to dividing the variance by the batch size.
@@ 4 Lower Bound for the Vanilla Stochastic Proximal Gradient Method
Proposition 1.
For any $K\geq 1$ and any (predefined) stepsize coefficients $\{M_{k}\}_{k=0}^{K-1}$ (possibly depending on the problem parameters $L,\sigma^{2}$ and $K$), there exists a problem instance of (1) with $f(\mathbf{x})\coloneqq\frac{L}{2}\left\lVert\mathbf{x}\right\rVert^{2}$ and $\psi(\mathbf{x})\coloneqq\frac{a}{2}\left\lVert\mathbf{x}\right\rVert^{2}$, $a\coloneqq\max_{0\leq k\leq K-1}M_{k}$, satisfying Assumptions 2 and 3, and the stochastic gradient oracle $\nabla f(\mathbf{x},\xi)\coloneqq\nabla f(\mathbf{x})+\xi$, $\xi\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$, satisfying Assumption 4, such that, for the sequence $\{\mathbf{x}_{k}\}_{k=1}^{K}$ generated by method (2) started from any initial point $\mathbf{x}_{0}$, it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k})\right\rVert^{2}\right]\geq\frac{1}{4}\sigma^{2}$ for any $1\leq k\leq K$.
@@
Clearly, the construction satisfies our assumptions. Let us fix an arbitrary $10\leq k\leq K-$ 1. Setting the gradient of the auxiliary function in (2) to zero, we see that $\mathbf{g}_{k}+a\mathbf{x}_{k+1}+M_{k}(\mathbf{x}_{k+1}-\mathbf{x}_{k})=0$. Since $\mathbf{g}_{k}=L\mathbf{x}_{k}+\xi_{k}$, it follows that
$\mathbf{x}_{k+1}=\frac{M_{k}-L}{M_{k}+a}\mathbf{x}_{k}-\frac{\xi_{k}}{M_{k}+a}.$
Substituting $\nabla F(\mathbf{x}_{k+1})=(L+a)\mathbf{x}_{k+1}$, we obtain
${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $=\frac{(L+a)^{2}}{(M_{k}+a)^{2}}{\mathbb{E}}\left[\left\lVert(M_{k}-L)\mathbf{x}_{k}-\xi_{k}\right\rVert^{2}\right]$
$=\frac{(L+a)^{2}}{(M_{k}+a)^{2}}\left((M_{k}-L)^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k}\right\rVert^{2}\right]+\sigma^{2}\right)$
$\geq\frac{(L+a)^{2}\sigma^{2}}{(M_{k}+a)^{2}}.$
If $a\geq M_{k}$, then $\frac{L+a}{M_{k}+a}\geq\frac{1}{2}$, and the claim follows. ∎
Note that Proposition 1 holds for any initial point $\mathbf{x}_{0}$. Even if we start at the optimal point $\mathbf{x}_{0}=\mathbf{0}$, the very first step of the method will already incur an $\mathcal{O}(\sigma^{2})$ error and no subsequent steps will be able to reduce it.
It is important that, for the composite optimization problem (1), we do not make any significant assumptions on $\psi$. In particular, we do not assume that $\psi$ is smooth with a certain smoothness constant. Proposition 1 demonstrates that, without such extra assumptions, for any fixed choice of parameters for the stochastic proximal gradient method (2), i.e., the stepsize coefficients $M_{k}$ and the number of iterations $K$, there is always a “bad” function in our problem class (namely, $F(\mathbf{x})=f(\mathbf{x})+\psi(\mathbf{x})$ with $f(\mathbf{x})=\frac{L}{2}\left\lVert\mathbf{x}\right\rVert^{2}$ and $\psi(\mathbf{x})=\frac{a}{2}\left\lVert\mathbf{x}\right\rVert^{2}$ for a sufficiently large $a$) for which the method cannot reach any error $<\frac{1}{4}\sigma^{2}$ after $K$ steps. In other words, for any given target accuracy $\varepsilon<\frac{1}{4}\sigma^{2}$, it is impossible to find one specific choice of the parameters for the method allowing it to reach the $\varepsilon$-error on any problem from our class.
The problem is that the variance of the gradient noise keeps the iterates away from the stationary point of $F$, and we need some mechanism reducing this variance with time. One such mechanism is the Polyak momentum which we discuss next.